OpenStudy (anonymous):

Get rid of the fraction by multiplying by the common denominator (x-2)/18 = (x-3)/15

OpenStudy (anonymous):

270

OpenStudy (anonymous):

15x-30+18x-54 on top? @mathway

OpenStudy (anonymous):

but i have a common denominator of three

OpenStudy (anonymous):

Are you finding for x?

OpenStudy (anonymous):

it just says get rid of the fractions by finding the common denominator on the paper

OpenStudy (anonymous):

The common denominator is 3.

OpenStudy (anonymous):

Yeah, after I find the common denominator then what do I do to the top? Does it change?

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

No

OpenStudy (anonymous):

270 is actually the common deoninator

OpenStudy (anonymous):

ugghhh i'm confused

OpenStudy (anonymous):

270 works because it's 18*15. Any number that both 18 and 15 go into works. I think 90 is the smallest. I think this is what you're trying to do\[\frac{ 270(x-2) }{ 18 }=\frac{ 270(x-3) }{ 15 }\] and then reduce the whole numbers \[15(x-2)=18(x-3)\] Which is exactly what you got when you cross multiplied

OpenStudy (anonymous):

see you made me confused a while ago >.< thanks @peachpi !

OpenStudy (anonymous):

np